OpenStudy (anonymous):

Please help! will fan and medal!!!! Tina is making rectangular signs. Each poster has an area of 1800 square inches and the length is 30 inches more than the width. Find the length and width of her sign. If x represents the width of each sign, select all the equations that could be used to solve this problem. 2(x)+2(x+30)=1800 (x+30)(x)=1800 x(30x)=1800 x(x+30)=1800

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

Part 2 Find the dimensions of each rectangular sign. width = inches length = inches

OpenStudy (mathstudent55):

For the posted problem at the top, B is an answer, but it's not the only answer. Another equation is the same equation with just the commutative property used.

OpenStudy (mathstudent55):

For Part 2, you need to solve the equation in choice B or the other correct choice from Part 1. \((x+30)(x)=1800\) Distribute the x. \(x^2 + 30x = 1800\) Subtract 1800 from both sides. \(x^2 + 30x - 1800 = 0\) Now you can try to factor (hint: it's factorable) or use the quadratic formula.